Please help me solve this question :
The sum of two numbers is 97. The larger number is one more than three times the smaller number. Find the numbers.

Answer by dabanfield(803)

Please help me solve this question :
The sum of two numbers is 97. The larger number is one more than three times the smaller number. Find the numbers.
Let x be the smaller number and y the larger.
Then we are given;
1.) x + y = 97
2.) y = 3*x + 1
From 1.) we know that y = 97 - x so we can "substitute" 97 - x for y in equation 2.):
97 - x = 3*x + 1
Solve the equation above for x and then calculate y = 97 - x.